OpenStudy (anonymous):

Q(x)=x^5-3x^4+3x^3+x^2-4x+2 Find all of the zeros of Q, real and complex and state the multiplicities

OpenStudy (anonymous):

1 is a root!

OpenStudy (anonymous):

-1 too!

OpenStudy (anonymous):

so (x+1)(x-1) can be pulled out

OpenStudy (anonymous):

but what is a multiplictias?

OpenStudy (anonymous):

sec I have to go now for a bit, will be back and try to solve it for the municipalities you need to factor it!

OpenStudy (anonymous):

i will be on it in a sec

OpenStudy (anonymous):

(x^2-1)(x^3-3x^2+4x-2)

OpenStudy (anonymous):

x^3-3x^2+4x-2 1 is a root again so x^3-3x^2+4x-2=(x-1)(x^2-2x+2)

OpenStudy (anonymous):

x^2-2x+2 has roots at 1+i and 1-i

OpenStudy (anonymous):

so roots are 1,-1,1+i,1-i all with 1 multi apart from x=1 that has 2 multi

OpenStudy (anonymous):

You have benn a great help, thanks!